OpenStudy (yoyo2443):

What is the equation in standard form of a parallel line that passes through (0,-2)? The line parallel to (0,-2) is y=-4x+4

OpenStudy (lord_box):

Try using the same process I showed you, except using the fact that parallel lines have identical slopes.

OpenStudy (lord_box):

Point-Slope form --> y = mx + b --> move mx to the left

OpenStudy (yoyo2443):

\[y+2=-4(x-0)\]

OpenStudy (yoyo2443):

-4 being the slope cause of the y=-4x+4

OpenStudy (lord_box):

that's right, now to y = mx + b form

OpenStudy (yoyo2443):

\[y=-4x+2\]

OpenStudy (lord_box):

Careful, you subtracted 2 from the left side, so you should subtract from the right as well.

OpenStudy (yoyo2443):

But I did subtract... ? from \[y+2=-4(x-0)\] \[y=-4x+2\]

OpenStudy (lord_box):

\[y + 2 = -4x\]\[y + 2 - 2 = -4x - 2\]

OpenStudy (yoyo2443):

Oh right I forgot about the 0..

OpenStudy (yoyo2443):

\[y+2-2=-4x-2\] \[y=-4x-2\] \[4x+y=-2\]?

OpenStudy (skullpatrol):

$$4x + y = -2$$ is correct :-)

OpenStudy (yoyo2443):

Thanks you two!

OpenStudy (skullpatrol):

Thanks for asking.